### v2

# references.bib

Abstract
The computational complexity of polynomial ideals and Gröbner bases has been studied since the 1980s. In recent years, the related notions of polynomial subalgebras and SAGBI bases have gained more and more attention in computational algebra, with a view towards effective algorithms. We investigate the computational complexity of the subalgebra membership problem and degree bounds. In particular, we show completeness for the complexity class EXPSPACE and prove PSPACE-completeness for homogeneous algebras. We highlight parallels and differences compared to the settings of ideals, and also look at important classes of polynomials such as monomial algebras.
1 Introduction
Let $\mathbb{K}$ be a field equipped with a suitable encoding over a finite alphabet, for example, the rational numbers $\mathbb{Q}$ or a finite field $\mathbb{F}_{q}$. At the heart of many algorithms in computer algebra lie efficient algorithms manipulating polynomial ideals and Gröbner bases, dating back at least to the 1960s [Buchberger2006]. Since the late 1980s, algorithms for computations with subalgebras have been studied [Shannon1988, Robbiano1990], which have applications in toric degenerations and polynomial system solving [betti2025]. A fundamental computational problem for ideals and subalgebras of polynomial rings $\mathbb{K}[{\underline{x}}]=\mathbb{K}[x_{1},\dots,x_{n}]$ is deciding membership. Of special interest are classes C of polynomials such as homogeneous polynomials (Homog), monomials (Mon), or polynomials with bounded degree or number of variables.
$\textsc{IdealMem}_{\mathbb{K}}(\textsf{C})$, $\textsc{IdealMem}_{\mathbb{K}}$
Input: $f_{1},\dots,f_{s}\in\textsf{C}$ (or $\mathbb{K}[{\underline{x}}]$);
$g\in\mathbb{K}[{\underline{x}}]$
Output: Is $g\in\langle f_{1},\dots,f_{s}\rangle$?
$\textsc{AlgMem}_{\mathbb{K}}(\textsf{C})$, $\textsc{AlgMem}_{\mathbb{K}}$
Input: $f_{1},\dots,f_{s}\in\textsf{C}$ (or $\mathbb{K}[{\underline{x}}]$);
$g\in\mathbb{K}[{\underline{x}}]$
Output: Is $g\in\mathbb{K}[f_{1},\dots,f_{s}]$?
Recall that $g\in\langle f_{1},\dots,f_{s}\rangle$ if and only if there is a representation $g=h_{1}f_{1}+\dots+h_{s}f_{s}$ with $h_{1},\dots,h_{s}\in\mathbb{K}[{\underline{x}}]$; the representation degree (of the tuple $f_{1},\dots,f_{s},g$) is the smallest possible value of $D=\max\{\deg h_{1},\dots,\deg h_{s}\}$. Similarly, $g\in\mathbb{K}[f_{1},\dots,f_{s}]$ if and only if $g=p(f_{1},\dots,f_{s})$ for some certificate $p\in\mathbb{K}[t_{1},\dots,t_{s}]$; we call the smallest possible degree of $p$ the certification degree.
The computational complexity and representation degree bounds of $\textsc{IdealMem}_{\mathbb{K}}$ and its variants have been studied since the 1980s, we give a brief overview in Section 2.3. The complexity of algebras in a more general sense has been studied before [Kozen77], though so far not for the concrete case of polynomial subalgebras. This paper is concerned with filling this gap and proving analogous results for $\textsc{AlgMem}_{\mathbb{K}}$ and bounds on the certification degree. In order to obtain reasonable upper bounds on the computational complexity, we assume that arithmetic in the base field $\mathbb{K}$ can be performed efficiently; this is formalized as being well-endowed [Borodin1983]. A main theorem is the EXPSPACE- and PSPACE-completeness of the subalgebra membership problem and the homogeneous variant, combining the upper bounds Theorem 3.1 and 3.4 and the complementing lower bound Theorem 4.2.
Theorem 1.1.
Over a well-endowed field $\mathbb{K}$, the subalgebra membership problem $\emph{{AlgMem}}_{\mathbb{K}}$ is EXPSPACE-complete and the homogeneous version $\emph{{AlgMem}}_{\mathbb{K}}(\textsf{{Homog}})$ is PSPACE-complete.
One of the goals of this paper is to be accessible both to commutative algebraists and computer scientists. For this reason, we introduce the relevant results from algebra and complexity theory on ideals and Gröbner bases in Section 2. In Section 3 we prove various upper bound results on variations of $\textsc{AlgMem}_{\mathbb{K}}$, as well as an upper bound on the certification degree. In Section 4 we present a complexity-theoretic reduction from $\textsc{IdealMem}_{\mathbb{K}}$ to $\textsc{AlgMem}_{\mathbb{K}}$, proving matching lower bounds. In Section 5 we study the case of monomial subalgebras, which is still NP-complete, and discuss related questions with SAGBI bases.
2 Background in algebra and complexity theory
2.1 Ideals, subalgebras and their bases
Let $\mathbb{K}$ be a field, denote by $\mathbb{K}[{\underline{x}}]\coloneqq\mathbb{K}[x_{1},\dots,x_{n}]$ the polynomial ring in $n$ variables, consisting of finite linear combinations
$f=\sum_{\alpha\in\mathbb{N}^{n}}c_{\alpha}x^{\alpha},\qquad x^{\alpha}=x_{1}^{\alpha_{1}}\dotsm x_{n}^{\alpha_{n}},\quad c_{\alpha}\in\mathbb{K}.$
Ideals and subalgebras arise naturally in algebra as the kernels and images of ($\mathbb{K}$-linear) ring homomorphisms. We briefly recall the basic definitions for both of them in parallel.
Definition 2.1.
A subset $I\subseteq\mathbb{K}[{\underline{x}}]$ is an ideal if it is an additive subgroup (contains 00 and is closed under addition and subtraction) and also closed under multiplication with arbitrary polynomials $f\in\mathbb{K}[{\underline{x}}]$. A subset $A\subseteq\mathbb{K}[{\underline{x}}]$ is a subalgebra if it is an additive subgroup containing $\mathbb{K}$ and is closed under multiplication within $A$. Given a set $S\subseteq\mathbb{K}[{\underline{x}}]$, we denote by
$\langle S\rangle$ =$=\Set{\sum_{\text{finite}}h_{i}f_{i}}{\!\!\!\begin{array}[]{c}f_{i}\in S,\\ h_{i}\in\mathbb{K}[{\underline{x}}]\end{array}\!\!\!}=\set{\text{lin.\ comb.\ of $S$ over $\mathbb{K}[{\underline{x}}]$}}$
$\mathbb{K}[S]$ =$=\Set{\sum_{\text{finite}}c_{\alpha}f^{\alpha}}{\!\!\!\begin{array}[]{c}f_{i}\in S,\\ c_{\alpha}\in\mathbb{K}\end{array}\!\!\!}=\set{\text{polynomial expressions in$S$}}$
the ideal and subalgebra generated by $S$; it is the smallest ideal resp. subalgebra containing $S$.
Example 2.2.
An important class of ideals and algebras are those generated by monomials $x^{\alpha}$, or binomials $x^{\alpha}-x^{\beta}$. For example, monomial algebras show up as coordinate rings of affine toric varieties, and binomial ideals are related to commutative Thue systems, capturing the combinatorial worst-case complexity of ideal membership (see Theorem 2.12).
Definition 2.3.
A monomial order $\prec$ is a total order on the set of monomials $x^{\alpha}$ stable under multiplication ($x^{\alpha}\preceq x^{\beta}$ implies $x^{\alpha}x^{\gamma}\preceq x^{\beta}x^{\gamma}$) and such that $1=x^{0}$ is the minimal element. The initial term $\operatorname{in}_{\prec}(f)$ of a nonzero polynomial $f=\sum_{\alpha}c_{\alpha}x^{\alpha}$ is the nonzero term $c_{\alpha}x^{\alpha}$ with the largest monomial. For convenience define $\operatorname{in}_{\prec}(0)=0$. Given an ideal $I$ or a subalgebra $A$, its initial ideal or initial algebra, respectively, is
$\operatorname{in}_{\prec}(I)=\langle\set{\operatorname{in}_{\prec}(f)}{f\in I}\rangle,\quad\operatorname{in}_{\prec}(A)=\mathbb{K}[\set{\operatorname{in}_{\prec}(f)}{f\in A}].$
In the following fix such a monomial order, for example, the lexicographic order $\prec_{\text{lex}}$; we may assume $x_{1}>\dots>x_{n}$.
Definition 2.4.
A Gröbner basis of an ideal $I$ is a finite subset $G\subseteq I$ such that $\operatorname{in}_{\prec}(I)=\langle\set{\operatorname{in}_{\prec}(g)}{g\in G}\rangle$. A Gröbner basis is reduced if the leading coefficients are $1$ and no term of $g\in G$ is in $\operatorname{in}_{\prec}(G\setminus\{g\})$.
On the other hand, a SAGBI basis of a subalgebra $A$ is any subset $S\subseteq A$ such that $\operatorname{in}_{\prec}(A)=\mathbb{K}[\set{\operatorname{in}_{\prec}(f)}{f\in S}]$.
In both cases, such bases generate $I$ and $A$ respectively, hence the (historical) name “basis”. The reduced Gröbner basis of an ideal $I$ is unique and has the smallest number of elements and smallest degree among all Gröbner bases of $I$. For more background on Gröbner bases see for example [Cox2016, Chapter 2], [Gathen2017, Chapter 21] or [Kreuzer2000, Chapter 2].
By Hilbert’s basis theorem, every ideal can be generated by a finite set of polynomials; in fact one can find such a finite set in any generating set. This implies the existence of (finite) Gröbner bases for any ideal. On the other hand, finite SAGBI bases may not exist; we will come back to this topic in Section 5.1.
Definition 2.5.
Given an ideal $I\subseteq\mathbb{K}[{\underline{x}}]$, the normal form $\operatorname{nf}_{\prec}(f,I)$ of an element $f\in\mathbb{K}[{\underline{x}}]$ against $I$ is the unique polynomial $r\in f+I$ such that no monomial of $r$ is in $\operatorname{in}_{\prec}(I)$ [Kreuzer2000, Definition 2.4.8].
One can compute $\operatorname{nf}_{\prec}(f,I)$ by dividing $f$ by a Gröbner basis $G$, the remainder is the normal form [Cox2016, Proposition 2.6.1]. This gives a membership test for polynomial ideals: $f\in I$ if and only if $\operatorname{nf}_{\prec}(f,I)=0$.
2.2 Complexity theory
We use the Turing model of computation, though the details of this formalism are not required to follow this paper. Here we introduce the concepts and computational problems used in the sequel. A classical reference for computational complexity theory is the book by Hopcroft and Ullman [Hopcroft1979], a more modern treatment is the book by Arora and Barak [Arora2009], for a brief introduction with a view towards computer algebra, see also [Gathen2017, Section 25.8].
Informally, the complexity of an algorithm is the amount of resources (time or memory) used in the computation as a function of the input length.
Definition 2.6.
A decision problem $A$ is the problem of deciding whether a given input $x\in\Sigma^{*}\coloneqq\{\text{words over the alphabet }\Sigma\}$ has a certain property, i. e. deciding membership in the set $A\subseteq\Sigma^{*}$. Problems in P, PSPACE, and EXPSPACE can be solved algorithmically in polynomial time, polynomial space, and exponential space, respectively, while the answer to problems in NP can be verified in polynomial time provided a certificate.
One has the inclusions of complexity classes
[…]
A standard problem in NP is Sat, deciding whether a given boolean formula $\varphi(x_{1},\dots,x_{n})$ can be satisfied by some assignment $\{x_{1},\dots,x_{n}\}\to\{\texttt{true},\texttt{false}\}$. We will later need the useful variant 1in3Sat (compare [Schaefer78], also known as positive/monotone 1in3Sat) and the UnboundedSubsetSum problem:
1in3Sat
Input: $S_{1},\dots,S_{n}\subseteq\mathbb{N}$, $|S_{i}|\leq 3$
Output: Is there a set $T\subseteq\mathbb{N}$ such that $|T\cap S_{i}|=1$ for all $i$?
UnboundedSubsetSum
Input: $a_{1},\dots,a_{s};b\in\mathbb{N}$
Output: Is $\sum_{i=1}^{s}c_{i}a_{i}=b$ for some $c_{1},\dots,c_{s}\in\mathbb{N}$?
A typical example of a problem in PSPACE is the halting problem for (deterministic) linearly bounded automata (LBA). Here a LBA $M$ consists of a finite set of states $Q$ including a starting state $q_{0}$ and a halting state $q_{\text{halt}}$, a tape alphabet $\Gamma=\{\texttt{0},\texttt{1},{\triangleright},{\triangleleft}\}$ containing the input alphabet $\Sigma=\{\texttt{0},\texttt{1}\}$, and a transition function
$\delta\colon(Q\setminus\{q_{\text{halt}}\})\times\Gamma\to Q\times\Gamma\times\{{\rm L},{\rm R}\}.$
A configuration is a triple $(q,i,b_{0}b_{1}\dots)\in Q\times\mathbb{Z}\times\Gamma^{*}$, and if $\delta(q,b_{i})=(q^{\prime},c,X)$, then $M$ transitions as
$(q,i,\dots b_{i-1}b_{i}b_{i+1}\dots)\Rightarrow\begin{cases}(q^{\prime},i-1,\dots b_{i-1}cb_{i+1}\dots)&\text{if }X={\rm L},\\ (q^{\prime},i+1,\dots b_{i-1}cb_{i+1}\dots)&\text{if }X={\rm R}.\end{cases}$
On the input string $w\in\Sigma^{*}$, the machine $M$ starts on the starting configuration $(q_{0},1,{\triangleright}w_{1}\dots w_{n}{\triangleleft})$ (here $\triangleright$ is at position 00) and repeatedly transitions according to $\delta$. The end markers $\triangleright$,$\triangleleft$ , ◁ may never be overwritten and the head may not pass over them (e.g. $\delta(q,{\triangleleft})=(q^{\prime},{\triangleleft},{\rm R})$ is not allowed). $M$ halts if it eventually reaches a configuration of the form $(q_{\text{halt}},i,b)$.
LBAhalt
Input: A LBA $M=(Q,\delta,q_{0},q_{\text{halt}})$ and an input $w\in\{\texttt{0},\texttt{1}\}^{*}$
Output: Does $M$ halt on input $w$?
Example 2.7.
The following table shows the transition function of a LBA consists of two states $q_{0},q_{1}$ (plus a halting state), and on input $w\coloneqq\texttt{0}\dots\texttt{0}$ ($n$ zeros) will count in binary to $\texttt{1}\dots\texttt{1}$ and then halt:
$\begin{array}[]{c|cccc}b&{\triangleright}&\texttt{0}&\texttt{1}&{\triangleleft}\\ \hline\cr\hline\cr q_{0}&&(q_{1},\texttt{1},{\rm L})&(q_{0},\texttt{0},{\rm R})&\text{halt}\\ q_{1}&(q_{0},{\triangleright},{\rm R})&(q_{1},\texttt{0},{\rm L})&&\end{array}$
Transitions that do not occur (on input $w$) have been left unspecified for simplicity. An example of the transition sequence can be found at
mathrepo.mis.mpg.de/ComplexityOfSubalgebras/configurations.html
Definition 2.8.
A complexity upper bound for a decision problem $A$ and a complexity class C is an algorithm solving $A$ and satisfying the resource constraints of C.
A lower bound or C-hardness result for $A$ is a proof that the computation of any problem $A^{\prime}\in\textsf{{C}}$ can be reduced to the problem $A$. We use the notion of log-space many-one reductions:
Definition 2.9.
For two sets $A\subseteq\Sigma^{*}$, $A^{\prime}\subseteq(\Sigma^{\prime})^{*}$, the problem $A^{\prime}$ is log-space many-one reducible to $A$, in symbols $A^{\prime}\mathrel{\leq^{\textsf{{L}}}_{\textup{m}}}A$, if there is a map $f\colon(\Sigma^{\prime})^{*}\to\Sigma^{*}$ computable in logarithmic working space, such that $x\in A^{\prime}$ if and only if $f(x)\in A$.
A problem $A$ is C-hard, if $A^{\prime}\mathrel{\leq^{\textsf{{L}}}_{\textup{m}}}A$ for all $A^{\prime}$ in C. It is C-complete if it is also in C.
For example, Sat, 1in3Sat, and UnboundedSubsetSum are NP-complete, and LBAhalt is PSPACE-complete.
In the problems $\textsc{IdealMem}_{\mathbb{K}}$ and $\textsc{AlgMem}_{\mathbb{K}}$, the input consists of encoded polynomials over $\mathbb{K}$, for example as a string of characters. Choosing a reasonable representation, the input length is bounded below by the number of terms, the number of variables occurring, and, depending on the encoding, the coefficients and the (logarithm of the) degree.
Definition 2.10.
The encoding of a polynomial $f=\sum_{|\alpha|\leq d}c_{\alpha}x^{\alpha}$ is dense if it lists all monomials with their coefficients $(\alpha,c_{\alpha})$ up to its degree, and sparse if it only lists those terms with non-zero coeffient. The representation of a monomial $x^{\alpha}\mathbin{\hat{=}}(\alpha_{1},\dots,\alpha_{n})\in\mathbb{N}^{n}$ is in binary if each $\alpha_{i}$ is encoded in binary, and unary if it is encoded in unary.
Example 2.11.
The term $f=42x_{1}^{3}x_{3}^{6}\in\mathbb{Q}[x_{1},x_{2},x_{3}]$ could be encoded with binary exponents as $((11,0,110),\texttt{enc}(42))$, and with unary as $((111,0,111111),\texttt{enc}(42))$. In dense encoding, a representation of $f$ would have to list all $\binom{9+3}{3}=220$ monomials up to degree $9$ (which most have coefficient 00), while a sparse encoding only has to list that one term.
All our complexity results will hold with respect to both dense and sparse encodings and binary or unary monomial representation (with the exception of Theorem 5.2), as long as we make the reasonable assumption that the field $\mathbb{K}$ is well-endowed [Borodin1983]. Commonly used fields in computer algebra such as number fields and finite fields are well-endowed. For this reason, we will not elaborate further on the encoding.
2.3 The ideal world: brief summary of results
In this section, we give an overview of the complexity results regarding ideal membership, representation degree, and Gröbner basis degrees. Our upper bounds on $\textsc{AlgMem}_{\mathbb{K}}$ will rely on refined representation and Gröbner basis degree upper bounds presented below. A more comprehensive discussion of the complexity of ideals can be found in [Mayr2017].
Theorem 2.12.
Let $\mathbb{K}$ be a well-endowed field (the lower bounds do not require this).
  1. (i)
    (Mayr & Mayer [Mayr1982], Mayr [Mayr1989]) The problem $\emph{{IdealMem}}_{\mathbb{K}}$ is complete for EXPSPACE. A representation can be enumerated in exponential working space.
  2. (ii)
    (Mayr [Mayr1997]) The problem $\emph{{IdealMem}}_{\mathbb{K}}(\textsf{{Homog}})$ is complete for PSPACE. The general problem is also in PSPACE when bounding the number of variables.
All hardness results already hold for ideals generated by binomials $x^{\alpha}-x^{\beta}$.
Remark 2.13.
Note that this uses the convention that in a computational problem with output, we only consider the working space and not the space needed to store the entire output. For example, one can enumerate the binary numbers $10,1,\dots,2^{n}-$ 1 with only $\mathcal{O}(n)$ bits of working memory. Similarly, here the number of terms in a representation $(h_{i})_{i}$ may be doubly exponential in the input length; this is unavoidable in the worst case [Mayr1982].
Representation degree bounds are, in general, doubly exponential in the number of variables, which is asymptotically optimal. We also note the special case of complete intersections for later reference. For a definition of complete intersections see e. g. [Cox2016, Exercise 9.4.8] or [Kreuzer2000, Definition 3.2.23].
Theorem 2.14.
Let $g\in\langle f_{1},\dots,f_{s}\rangle\subseteq\mathbb{K}[x_{1},\dots,x_{n}]$, $d\coloneqq\max_{i}\deg f_{i}$, and write $g=\sum_{i=1}^{s}h_{i}f_{i}$ with $h_{i}\in K[{\underline{x}}]$ of minimal representation degree $D\coloneqq\max_{i}\deg h_{i}$.
  1. (i)
    (Hermann [Hermann1926], Mayr & Meyer [Mayr1982]) $D\leq\deg g+(ds)^{2^{n}}$.
  2. (ii)
    (Dickenstein, Fitchas, Giusti & Sessa [Dickenstein1991]) If $f_{1},\dots,f_{s}$ define a complete intersection ideal, then $D\leq\deg g+d^{s}$.
Both from a theoretical and computational point of view, effective degree bounds on Gröbner bases are important as well. We recall the classical Dubé bound as well as a dimension-refined version. Here, the dimension of an ideal $I\subseteq\mathbb{K}[{\underline{x}}]$ is the (Krull) dimension of the quotient ring $\mathbb{K}[{\underline{x}}]/I$, see [Cox2016, Chapter 9] for other characterizations of the dimension.
Theorem 2.15.
Let $I\subseteq\mathbb{K}[x_{1},\dots,x_{n}]$ be an ideal generated by $f_{1},\dots,f_{s}$ of degree $d_{i}\coloneqq\deg f_{i}$, $d_{1}\geq\dots\geq d_{s}$, and consider any monomial order.
  1. (i)
    (Dubé [Dube1990]) The degree of the reduced Gröbner basis GB of $I$, i. e. the largest degree of an element of GB, is bounded by
    $\deg\textit{GB}\leq 2\left(\frac{1}{2}d_{1}^{2}+d_{1}\right)^{2^{n-1}}.$
  2. (ii)
    (Mayr & Ritscher [Mayr2013]) If $I$ has dimension $r$, then the degree is bounded by
    $\deg\textit{GB}\leq 2\left(\frac{1}{2}(d_{1}\dotsm d_{n-r})^{2(n-r)}+d_{1}\right)^{2^{r}}\!\leq 2\left(\frac{1}{2}d_{1}^{2(n-r)^{2}}+d_{1}\right)^{2^{r}}.$
These upper bounds are asymptotically sharp.
3 Upper bounds on subalgebra membership
In this section we provide various complexity upper bounds, that is, algorithms for variants of $\textsc{AlgMem}_{\mathbb{K}}$ which place these problems in PSPACE and EXPSPACE. Our complexity analysis of the membership problem for a subalgebra $A=\mathbb{K}[f_{1},\dots,f_{s}]\subseteq\mathbb{K}[{\underline{x}}]$ relies on the following classical elimination approach using tag variables, attributed to Spear [Shannon1988, Spear1977].
Let ${\underline{t}}=\{t_{1},\dots,t_{s}\}$ be additional variables, and consider an elimination order on $\mathbb{K}[{\underline{x}},{\underline{t}}]$, i. e. a monomial order such that $x_{i}\succ t^{\alpha}$ for all $x_{i}$ and $t^{\alpha}\in\mathbb{K}[{\underline{t}}]$, for example the lexicographical order $\prec_{\text{lex}}$.
{lemma}
[Shannon & Sweedler [Shannon1988]] Let $f_{1},\dots,f_{s},g\in\mathbb{K}[{\underline{x}}]$, $J\coloneqq\langle f_{1}-t_{1},\dots,f_{s}-t_{s}\rangle$, then
$g\in\mathbb{K}[f_{1},\dots,f_{s}]\qquad\text{if and only if}\qquad p\coloneqq\operatorname{nf}_{\prec}(g,J)\in\mathbb{K}[{\underline{t}}].$
If this is the case, then $g=p(f_{1},\dots,f_{s})$, interpreting $p$ as a polynomial in $t_{1},\dots,t_{s}$.
Subalgebra membership can thus be reduced to the task of calculating the normal form of a polynomial against an ideal in a larger polynomial ring. This approach is well-known, for example as a fall-back method in the Macaulay2 [M2] package SubalgebraBases [Burr2024]. Kühnle & Mayr [Kühnle1996] describe an exponential space algorithm for enumerating the normal form of a polynomial over a well-endowed field. Combining these results gives the first complexity upper bound:
Theorem 3.1.
For any well-endowed field $\mathbb{K}$, $\emph{{AlgMem}}_{\mathbb{K}}$ is in EXPSPACE. Moreover a certificate $p\in\mathbb{K}[t_{1},\dots,t_{s}]$ such that $g=p(f_{1},\dots,f_{s})$ can be output in exponential working space.
Proof.
Given $f_{1},\dots,f_{s},g$, the algorithm computes the normal form $p\coloneqq\operatorname{nf}_{\prec}(g,J)\in\mathbb{K}[{\underline{x}},{\underline{t}}]$ using Kühnle & Mayr’s algorithm [Kühnle1996]. If there is a nonzero term in $p$ involving some $x_{i}$, then $g\notin\mathbb{K}[f_{1},\dots,f_{s}]$. Otherwise, $p$ is a certificate of subalgebra membership by Section 3, which can be enumerated to the output tape using single exponential working space.
Note that it might not be possible to fit the normal form in exponential working space as mentioned in Remark 2.13. Instead one has to enumerate the terms of $p$, which fit in exponential working space using and check for occurrences of $x_{i}$ individually. ∎
To prove degree bounds and better complexity upper bounds for special classes of polynomials, we outline part of Kühnle & Mayr’s upper bound constructions on normal forms [Kühnle1996].
[…]
In our specialized setting $J=\langle t_{1}-f_{1},\dots,t_{s}-f_{s}\rangle$ we have more refined upper bounds available regarding representation and Gröbner basis degrees. We apply this to the case of a fixed number of variables $n$ and to the case of a fixed subalgebra $A$.
Theorem 3.2.
For a fixed subalgebra $A=\mathbb{K}[f_{1},\dots,f_{s}]\subseteq\mathbb{K}[{\underline{x}}]$ over a well-endowed field, the membership problem is in PSPACE (with respect to the input length of $g$). In fact, for a fixed number of variables $n$, we have $\emph{{AlgMem}}_{\mathbb{K}}(\mathbb{K}[x_{1},\dots,x_{n}])\in\textsf{{PSPACE}}$.
Proof.
The algorithm is the same as in Theorem 3.1, only the complexity analysis is slightly more elaborate. Using that normal form calculation with respect to a fixed ideal is possible in polynomial space, as remarked by Kühnle & Mayr [Kühnle1996, Section 6], the first assertion follows.
If only the number of variables $x_{1},\dots,x_{n}$ is fixed, then more care has to be taken, as the computation takes place in the ring $\mathbb{K}[x_{1},\dots,x_{n},t_{1},\dots,t_{s}]$, whose number of variables $n+s$ still depends on the input length. However, the ideal $J=\langle t_{1}-f_{1},\dots,t_{s}-f_{s}\rangle$ is a complete intersection of dimension $n$, hence using Theorem 2.15 its Gröbner basis degree is bounded above by
$\deg\textit{GB}(J)\leq G\coloneqq 2\left(\frac{1}{2}d^{2s^{2}}+d\right)^{2^{n}}.$
Furthermore, the representation degree of $g-\operatorname{nf}_{\prec}(g,J)$ is bounded above by $D\coloneqq G+d^{s}$ by Theorem 2.14. Since $n$ is fixed, we see that $D$ is single exponential in the input length, hence $\operatorname{polylog}(D)$ is polynomial. This shows that the Meyer & Kühnle algorithm works in polynomial space for a bounded number of variables. ∎
{corollary}
If $g\in\mathbb{K}[f_{1},\dots,f_{s}]\subseteq\mathbb{K}[{\underline{x}}]$, then there exists a certificate $p\in\mathbb{K}[t_{1},\dots,t_{s}]$ with $g=p(f_{1},\dots,f_{s})$ of degree
$\deg p\leq\deg(g)+\left(\left(\textstyle\frac{1}{2}d^{2s^{2}}+d\right)^{2^{n}}+1\right)^{(n+s)^{2}+1}\deg(g)^{n+s}.$
Proof.
This is the degree bound on the normal form by Kühnle & Mayr evaluated for the ideal $J$ (note that the ambient polynomial ring has $n+s$ variables). ∎
Remark 3.3.
This degree can probably be improved by studying the derivation of the normal form degree bounds in this particular situation. We suspect that a “nicer” upper bound akin to the Hermann bound (Theorem 2.14) should hold.
Of great interest is the case of subalgebras generated by homogeneous polynomials. In this case, $A$ is graded compatibly with the standard grading of the ambient ring.
Theorem 3.4.
For any well-endowed field $\mathbb{K}$, $\emph{{AlgMem}}_{\mathbb{K}}(\textsf{{Homog}})$ is in PSPACE.
Proof.
As a consequence of the grading, we see that a certificate $p$ of lowest degree (if it exists) has degree at most $\deg g$. Thus either $\deg\operatorname{nf}_{\prec}(g,J)>\deg g$, in which case $g\notin K[f_{1},\dots,f_{s}]$, or $\deg\operatorname{nf}_{\prec}(g,J)\leq\deg g$ and $g\in\mathbb{K}[f_{1},\dots,f_{s}]$ if and only if $\operatorname{nf}_{\prec}(g,J)\in K[{\underline{t}}]$. This leads to a variation of the algorithm of Theorem 3.1 except we only attempt to compute the normal form up to degree $\deg g$. The complexity analysis is analogous to Theorem 3.2, noting that $\deg\operatorname{nf}_{\prec}(g,J)\leq\deg g$, hence a representation degree bound here is $D\coloneqq\deg g+d^{s}$. ∎
Remark 3.5.
Degree bounds on certificates for elimination problems have been studied before, for example by Galligo & Jelonek [Galligo2020elimination]. Their results are not necessarily applicable here, as our ideal $J$ has dimension $n$ in a ring of $n+s$ variables, while the results of Galligo & Jelonek only apply to intersections of an ideal with $K[x_{1},\dots,x_{n}]$ if the ideal has dimension at least $n+1$.
4 Lower bounds on subalgebra membership
In this section, we prove matching complexity lower bounds for the subalgebra membership problem and its homogeneous variant, relating them to the construction for homogeneous ideals. The hardness results for $\textsc{IdealMem}_{\mathbb{K}}$ and $\textsc{IdealMem}_{\mathbb{K}}(\textsf{{Homog}})$ were constructed by embedding the word problem for commutative semigroups resp. its homogeneous variant into binomial ideals [Mayr1982, Mayr1997]. We now describe how to embed (binomial) ideal membership into subalgebra membership.
Theorem 4.1.
Let $f_{1},\dots,f_{s},g\in\mathbb{K}[x_{1},\dots,x_{n}]$ and let $t$ be a new variable. The following are equivalent:
  1. (a)
    $g\in\langle f_{1},\dots,f_{s}\rangle\subseteq\mathbb{K}[{\underline{x}}]$;
  2. (b)
    $tg\in\mathbb{K}[tf_{1},\dots,tf_{s},x_{1},\dots,x_{s}]$.
Proof.
If $g=h_{1}f_{1}+\dots+h_{s}f_{s}$ with $h_{i}\in\mathbb{K}[{\underline{x}}]$, then
$tg=h_{1}\cdot(tf_{1})+\dots+h_{s}\cdot(tf_{s}),$
which certifies that $tg$ is in $\mathbb{K}[tf_{1},\dots,tf_{s},{\underline{x}}]$.
Conversely, assume $tg=p(tf_{1},\dots,tf_{s},x_{1},\dots,x_{n})$ with $p\in R\coloneqq\mathbb{K}[u_{1},\dots,u_{s},x_{1},\dots,x_{n}]$. Declare a (non-standard) grading on $R$ and $S=\mathbb{K}[t,x_{1},\dots,x_{n}]$ by giving $\deg t=\deg u_{j}=1$ and $\deg x_{i}=0$, so
$R_{d}\coloneqq\bigoplus_{\begin{subarray}{c}\alpha\in\mathbb{N}^{s}\\ |\alpha|=d\end{subarray}}u_{1}^{\alpha_{1}}\dotsm u_{s}^{\alpha_{s}}\mathbb{K}[x_{1},\dots,x_{n}],\qquad S_{d}\coloneqq t^{d}\mathbb{K}[x_{1},\dots,x_{n}].$
The evaluation homomorphism $R\to S$ substituting $x_{i}\mapsto x_{i}$, $u_{j}\mapsto t_{j}f_{j}$ respects this grading. Since $tg\in S_{1}$ is homogeneous of degree $1$, decomposing $p$ into its graded components $p=\sum_{d}p_{d}$ ($p_{d}\in R_{d}$) reveals that
$p_{d}(tf_{1},\dots,tf_{s},x_{1},\dots,x_{n})=\begin{cases}p(tf_{1},\dots,x_{n})=tg&\text{if }d=1\\ 0&\text{if }d\neq 1.\end{cases}$
Thus by replacing $p$ by $p_{1}$, we may assume that every term in $p$ involves exactly one $u_{j}$. Setting $t=1$, we read
$p(f_{1},\dots,f_{s},x_{1},\dots,x_{n})=g,$
which, by the structure of $p$, is a representation for $g\in\langle f_{1},\dots,f_{s}\rangle$. ∎
{corollary}
Follow the notation from the previous theorem and set $A\coloneqq\mathbb{K}[tf_{1},\dots,tf_{s},{\underline{x}}]$.
  1. (i)
    The representation degree of $g\in\langle f_{1},\dots,f_{s}\rangle$ is one less than the certification degree of $tg\in A$.
  2. (ii)
    The smallest possible number of non-zero terms of all $h_{i}$ in a representation $g=h_{1}f_{1}+\dots+h_{s}f_{s}$ equals the smallest possible number of terms of a certificate for $tg\in A$.
Proof.
The forward direction in the previous proof gives that the total degree and number of terms for representation of ideal membership are a lower bound for the respective values for the subalgebra (the total degree increases by $1$ by attaching the $u_{j}$ to the $h_{j}$). Conversely, the grading argument shows that if $p$ is a certificate, then so is its graded component $p_{1}$, which has a lower or equal degree and number of terms. Finally, setting $t\mapsto 1$ in $p_{1}$ gives a representation for $g\in\langle f_{1},\dots,f_{s}\rangle$, showing the desired equalities. ∎
With this result we can deduce matching complexity lower bounds for subalgebra membership and its homogeneous variant. We can rephrase Theorem 4.1 as a complexity-theoretic reduction.
Theorem 4.2.
The map
$(f_{1},\dots,f_{s};g\in\mathbb{K}[{\underline{x}}])\;\mapsto\;(tf_{1},\dots,tf_{s},x_{1},\dots,x_{n};tg\in\mathbb{K}[t,{\underline{x}}])$
provides a reduction $\emph{{IdealMem}}_{\mathbb{K}}\mathrel{\leq^{\textsf{{L}}}_{\textup{m}}}\emph{{AlgMem}}_{\mathbb{K}}$. More generally, it provides a reduction $\emph{{IdealMem}}_{\mathbb{K}}(\textup{{C}})\mathrel{\leq^{\textsf{{L}}}_{\textup{m}}}\emph{{AlgMem}}_{\mathbb{K}}(\textup{{C}})$ for any class of polynomials C containing single variables and closed under multiplication with variables.
With this we can complete the proof of our main theorem.
Proof of Theorem 1.1.
The complexity-theoretic upper bounds have been established in Theorem 3.1 and Theorem 3.4. The lower bounds now follow from the hardness of $\textsc{IdealMem}_{\mathbb{K}}$ and its homogeneous variant (Theorem 2.12) in conjunction with the reduction in Theorem 4.2. ∎
As mentioned in the beginning of this section, the complexity lower bounds for ideal membership were achieved by studying the word problem for commutative semigroups. Let =$\mathcal{R}=\{(x^{\alpha_{i}},x^{\beta_{i}})\}_{i=1}^{s}$ be a set of pairs of monomials, called replacement rules. Let $\equiv_{\mathcal{R}}$ be the congruence relation on $\textsf{{Mon}}({\underline{x}})$ generated by
$x^{\gamma}x^{\alpha}\equiv_{\mathcal{R}}x^{\gamma}x^{\beta}\qquad(x^{\alpha},x^{\beta})\in\mathcal{R},\ x^{\gamma}\in\textsf{{Mon}}({\underline{x}}).$
Concretely, two monomials $m,m^{\prime}$ are equivalent if one can be turned into the other by repeatedly replacing a factor $x^{\alpha}\mid m$ by the corresponding $x^{\beta}$ or by replacing $x^{\beta}\mid m$ by $x^{\alpha}$.
Word problem for Commutative Semigroups, CSG
Input: Replacement rules =$\mathcal{R}=\{(x^{\alpha_{i}},x^{\beta_{i}})\}_{i=1}^{s}$,
Monomials $m,m^{\prime}\in\textsf{{Mon}}({\underline{x}})$
Output: Is $m\equiv_{\mathcal{R}}m^{\prime}$?
Mayr & Meyer [Mayr1982] prove EXPSPACE-hardness of CSG and provide a reduction $\textsc{CSG}\mathrel{\leq^{\textsf{{L}}}_{\textup{m}}}\textsc{IdealMem}_{\mathbb{K}}(\textsf{Binom})$ via the map
$(\mathcal{R},m,m^{\prime})\;\mapsto\;(x^{\alpha_{1}}-x^{\beta_{1}},\dots,x^{\alpha_{s}}-x^{\beta_{s}};m-m^{\prime}).$
More precisely, the minimal number of terms in a representation for this ideal membership problem equals the minimal number of replacement steps necessary in showing the congruence $m\equiv_{\mathcal{R}}m^{\prime}$. Using the known double-exponential examples from [Mayr1982], we can prove worst-case certification degree lower bounds for subalgebra membership. For an accessible exposition of the double-exponential lower bounds compare [Bayer1988] or the author’s Master’s thesis [kayser2022groebnercomplexity].
Theorem 4.3.
For given $k$ there exist polynomials $f_{1},\dots,f_{s},g\in\mathbb{K}[x_{1},\dots,x_{n}]$ of degree $\leq 6$, $n=\mathcal{O}(k)$, $s=\mathcal{O}(k)$, such that $g\in\mathbb{K}[f_{1},\dots,f_{s}]$, but every certificate $p\in\mathbb{K}[t_{1},\dots,t_{s}]$ has degree at least $2^{2^{n}}$ and at least $2^{2^{n}}$ terms. These $f_{i}$ and $g$ can be chosen as binomials and monomials.
Proof.
The CSG instances from [Mayr1982] lead to ideal membership of $\mathcal{O}(k)$ binomials in $\mathcal{O}(k)$ variables of degree $\leq 5$ (to obtain this exact degree bound, use the slightly improved degree bound from [Bayer1988]) such that every representation has degree and number of terms bounded below by $2^{2^{n}}$. Applying the reduction from Theorem 4.2, we obtain an instance of $\textsc{AlgMem}_{\mathbb{K}}$, whose smallest certificate $p$ has the same number of terms and greater total degree by Section 4. ∎
Remark 4.4.
Following [Bayer1988], this construction can be easily adapted to prove lower bounds of the form $d^{2^{k}}$ at the expense of increasing the algebra generator degree to $d+\mathcal{O}(1)$. This asymptotically matches the double-exponential upper bounds from Section 3. It would be interesting to get more refined lower bounds in terms of other properties of the algebra. For example, what could analogous statements of Theorem 2.14 and Theorem 2.15 be?
While the worst-case examples for CSG are too complicated to recall here, we can give more details in the homogeneous case. Here the certification degree can not blow up, however, we can still produce $\textsc{AlgMem}_{\mathbb{K}}(\textsf{{Homog}})$ instances that require exponentially many terms in their certificates.
Theorem 4.5.
There exist a subalgebra $A=\mathbb{K}[f_{1},\dots,f_{5n}]\subseteq\mathbb{K}[x_{1},\dots,x_{3n+3}]$ generated by single variables and homogeneous binomials of degree $\leq 3$, and a binomial $g$ of degree $n+2$ with the following property: $g\in A$, but every polynomial $p\in\mathbb{K}[{\underline{t}}]$ with $g=p(f_{1},\dots,f_{5n})$ has at least $2^{n+1}$ terms.
Proof.
The construction is based on the binary-counting LBA from Example 2.7. In [Mayr1997, Theorem 17] Mayr describes how to reduce LBAhalt to commutative semigroups/ideals, proving PSPACE-hardness of $\textsc{IdealMem}_{\mathbb{K}}(\textsf{{Homog}})$. Applying our reduction (Theorem 4.2), we obtain a binary-counting subalgebra. The configuration $(q,i,b_{0},\dots,b_{n+1})$ of the LBA is represented by the monomial $qh_{i}x_{0,{\triangleright}}x_{1,b_{1}}\dotsm x_{n,b_{n}}x_{n+1,{\triangleleft}}$ and the replacement rules/binomial generators mimic the transition function $\delta$.
[…]
$\mathcal{T}=\{q_{0},q_{1}\}$$\mathbin{\dot{\cup}}\{h_{0},\dots,h_{n}\},\quad{\underline{x}}=\{x_{1,\texttt{0}},x_{1,\texttt{1}},\dots,x_{n,\texttt{0}},x_{n,\texttt{1}}\},$
$\mathcal{R}$$\coloneqq\set{q_{0}h_{i}x_{i,\texttt{0}}-q_{1}h_{i-1}x_{i,\texttt{1}}}{1\leq i\leq n}$
${}\;\cup\set{q_{0}h_{i}x_{i,\texttt{1}}-q_{0}h_{i+1}x_{i,\texttt{0}}}{1\leq i\leq n-1}$
${}\;\cup\set{q_{1}h_{i}x_{i,\texttt{0}}-q_{1}h_{i-1}x_{i,\texttt{0}}}{1\leq i\leq n}$
${}\;\cup\set{q_{1}h_{0}-q_{0}h_{1}},$
$A$ $\coloneqq\mathbb{K}[f_{1},\dots,f_{5n}]\coloneqq\mathbb{K}[\mathcal{T}\cup{\underline{x}}],$
$g$ $\coloneqq q_{0}h_{1}x_{1,\texttt{0}}\dotsm x_{n,\texttt{0}}-q_{0}h_{n}x_{1,\texttt{0}}\dotsm x_{n-1,\texttt{0}}x_{n,\texttt{1}}.$
Then $g\in A$ by construction, since the binary counter will go from $(q_{0},1,\triangleright\texttt{0}\dots\texttt{0}\triangleleft)$ to $(q_{0},1,\triangleright\texttt{1}\dots\texttt{1}\triangleleft)$ and then walk to the right, erasing the 1’s until it reaches the configuration $(q_{0},n,\triangleright\texttt{0}\dots\texttt{0}\texttt{1}\triangleleft)$. On the way it writes every number $0,\dots,2^{n-1}$ on the tape, taking at least $2$ steps each time, for a total of $\geq 2^{n+1}$ steps. By Section 4 any certificate for $g\in A$ must essentially contain this derivation, hence has at least $2^{n+1}$ terms. ∎
An implementation of the homogeneous binary-counting subalgebra in Macaulay2 can be found at mathrepo.mis.mpg.de/ComplexityOfSubalgebras.
5 Monomial subalgebras and SAGBI bases
In this final section, we consider the complexity of monomial algebra membership and consider some questions related to SAGBI bases. Monomial subalgebras $A$ are $\mathbb{N}^{n}$-graded in the sense that
$A=\bigoplus_{\begin{subarray}{c}\alpha\in\mathbb{N}^{n}\\ x^{\alpha}\in A\end{subarray}}\mathbb{K}x^{\alpha}\subseteq\mathbb{K}[{\underline{x}}].$
This has the useful consequence that a polynomial $\sum_{\alpha}c_{\alpha}x^{\alpha}$ is in $A$ if and only if every monomial $x^{\alpha}$, $c_{\alpha}\neq 0$, is in $A$. Furthermore, monomial algebras are related to linear programming and linear Diophantine equations [Robbiano1990, Remark 1.9], as
$x^{\beta}\in K[x^{\alpha_{1}},\dots,x^{\alpha_{s}}]\quad\Longleftrightarrow\quad\exists c\in\mathbb{N}^{s}\text{ s.t.\ }\beta=\sum_{i=1}^{s}c_{i}\alpha_{s}.$ (1)
Theorem 5.1.
For any $\mathbb{K}$ the problem $\emph{{AlgMem}}_{\mathbb{K}}(\textsf{{Mon}})$ is NP-complete. This is true even when restricting to square-free monomials.
Proof.
For NP membership one immediately reduces to the case where the polynomial to test $f$ is a monomial due to the $\mathbb{N}^{n}$-grading. In Equation 1 we have $c_{j}\leq\max_{i}\beta_{i}$, so the bit length of $c_{j}$ is bounded by the bit length of $\beta$. Hence, non-deterministically guessing $c$ yields a NP-algorithm.
For NP-hardness, one can reduce from the NP-complete problem 1in3Sat (Section 2.2). Indeed, given sets $S_{1},\dots,S_{n}\subseteq\{1,\dots,s\}$, then let $\alpha_{1},\dots,\alpha_{s}\in\{0,1\}^{n}$ be the integer vectors with $(\alpha_{i})_{j}=1$ when $i\in S_{j}$. Set $\beta=(1,\dots,1)\in\mathbb{N}^{n}$. Then Equation 1 encodes exactly the question for 1in3Sat, a solution corresponding to $T=\set{i}{c_{i}=1}$. In this construction, all monomials $x^{\alpha_{i}},x^{\beta}$ are square-free. ∎
We see that $\textsc{AlgMem}_{\mathbb{K}}(\textsf{{Mon}})$ is NP-complete even for polynomials of degree $\leq n$. The same is true if we instead bound the number of variables – if the exponents are encoded in binary.
Theorem 5.2.
The problem $\emph{{AlgMem}}_{\mathbb{K}}(\textsf{{Mon}}(x_{1},\dots,x_{n}))$ for fixed $n\geq 1$ is NP-complete for binary exponent encoding and in TC0 with unary encoding.
[…]
$\texttt{maj}(b_{1},\dots,b_{n})=\begin{cases}1&\text{if }\#\set{i}{b_{i}=1}\geq\frac{n}{2}\\ 0&\text{otherwise.}\end{cases}$
For more information on TC0 and circuit complexity classes, see [Vollmer1999, Chapter 4].
Proof.
Encoding the exponents as binary, the unary case $n=1$ is a direct translation of the UnboundedSubsetSum problem, which is NP-hard.
On the other hand, if the monomials are encoded in unary, then a generating-function approach as in [Kane2017] provides a family of circuits in TC0 deciding $\textsc{AlgMem}_{K}(\textsf{{Mon}}(x_{1},\dots,x_{n}))$. In more detail, Kane describes TC0 circuits deciding the unary vector-valued subset sum problem, which corresponds to solutions of Equation 1 with $c_{i}\in\{0,1\}$. We can reduce the case of arbitrary $c_{i}\in\mathbb{N}$ to the $\{0,1\}$ case by replacing every generator $x^{\alpha_{i}}$ by $x^{\alpha_{i}},x^{2\alpha_{i}},x^{4\alpha_{i}},\dots,x^{2^{\kappa}\alpha_{i}}$, $\kappa=\lfloor\log_{2}|\beta|\rfloor$. Since the numbers are encoded in unary, the input size is proportional to $C=|\beta|+\sum_{i=1}^{s}|\alpha_{i}|$, and this larger generating set has size $\mathcal{O}(C^{2})$.
Alternatively, one can apply a TC0-variant of Courcelle’s theorem to obtain TC0-membership, compare [Elberfeld2012, Theorem 13]. ∎
Remark 5.3.
The NP-hardness results from Theorem 5.1 and 5.2 are in stark contrast to the analogous results for monomial ideals: Monomial ideal membership is computationally trivial, as $x^{\beta}\in\langle x^{\alpha_{1}},\dots,x^{\alpha_{s}}\rangle$ if and only if component-wise $\beta\geq\alpha_{i}$ for some $i$.
5.1 Some remarks on SAGBI bases
We return to a general subalgebra $A=\mathbb{K}[f_{1},\dots,f_{s}]\subseteq\mathbb{K}[{\underline{x}}]$ equipped with a monomial order $\prec$. In Section 2.1 we defined the concept of SAGBI bases $S\subseteq A$ as subsets with $\mathbb{K}[\set{\operatorname{in}_{\prec}f}{f\in S}]=\operatorname{in}_{\prec}A$. SAGBI stands for “Subalgebra Analog to Gröbner Bases for Ideals” [Robbiano1990] and they are used in practice for deciding subalgebra membership [Burr2024]. Much of the theory of Gröbner bases is paralleled for SAGBI bases, such as the subduction algorithm deciding subalgebra membership, reminiscent of the division algorithm for Gröbner bases [Kreuzer2005-rb, Section 6.6].
The previous results in this section provide a modest complexity lower bound of deciding subalgebra membership in the presence of SAGBI bases. The author is not aware of an analogous result for ideal membership given a Gröbner basis.
{corollary}
The problem $\emph{{AlgMem}}_{K}$ is NP-hard, even if the input polynomials $f_{1},\dots,f_{s}$ form a finite SAGBI basis of $\mathbb{K}[f_{1},\dots,f_{s}]$.
Proof.
Subalgebras generated by a finite set $S$ of monomials have $S$ as a SAGBI basis. Hence Theorem 5.1 provides the NP lower bound. ∎
A major difference to the ideal world is that not every finitely generated subalgebra has a finite SAGBI basis.
Example 5.4.
The subalgebra $\mathbb{K}[x_{1},x_{1}x_{2}-x_{2}^{2},x_{1}x_{2}^{2}]\subseteq\mathbb{K}[x_{1},x_{2}]$ has the non-finitely generated initial algebra $\mathbb{K}[\set{x_{1}x_{2}^{k}}{k\geq 0}]$ for any monomial order with $x_{1}\succ x_{2}$ [Robbiano1990, Example 4.11].
We therefore propose the study of the following two decision problems related to the initial algebra.
$\textsc{SAGBIfinite}_{\mathbb{K}}(\textsf{C})$, SAGBIfinite
Input: $f_{1},\dots,f_{s}\in\textsf{C}$ (or $\mathbb{K}[{\underline{x}}]$)
Output: Does $\mathbb{K}[f_{1},\dots,f_{s}]$ have a finite SAGBI basis?
$\textsc{InAlgMem}_{\mathbb{K}}(\textsf{C})$, $\textsc{InAlgMem}_{\mathbb{K}}$
Input: $f_{1},\dots,f_{s}\in\textsf{C}$ (or $\mathbb{K}[{\underline{x}}]$);
$x^{\alpha}\in\textsf{{Mon}}({\underline{x}})$
Output: Is $x^{\alpha}\in\operatorname{in}_{\prec}\mathbb{K}[f_{1},\dots,f_{s}]$?
Robbiano & Sweedler showed that an algorithm somewhat analogous to Buchberger’s algorithm for Gröbner bases can be used to enumerate a SAGBI basis, which will terminate if (and only if) $A$ has a finite SAGBI basis. A procedure that returns yes if the output is yes, but never terminates if the output is no, is called a semi-algorithm, and a problem is semi-decidable or recursively enumerable if there is a semi-algorithm “solving” it.
Theorem 5.5 (Robbiano & Sweedler [Robbiano1990]).
$\emph{{InAlgMem}}_{\mathbb{K}}$ and $\emph{{SAGBIfinite}}_{\mathbb{K}}$ are semi-decidable (over a computable field).
We are not aware of any general better complexity bounds, or even just if these problems are computable at all (though a negative answer would be quite surprising). Future work will provide a more detailed study of the structure and complexity of (infinitely generated) initial algebras.
Acknowledgments
I would like to thank my advisor, Simon Telen, as well as Markus Bläser, Peter Bürgisser, Florian Chudigiewitsch, and Fulvio Gesmundo for helpful discussions, in particular Florian for suggesting Courcelle’s theorem for Theorem 5.2. I am also indebted to the anonymous reviewers, whose reports helped to improve the presentation and significantly simplify the treatment of Section 4. My interest in the complexity of ideals and subalgebras originated from my Master’s thesis [kayser2022groebnercomplexity] supervised by Heribert Vollmer & Sabrina Gaube at Leibniz University Hannover.
[…]
Author’s address:
Leonie Kayser, MPI-MiS Leipzig leo.kayser@mis.mpg.de